OpenStudy (anonymous):

Write your final answer exactly as a fraction in lowest terms. 22/45 divided 14/25

OpenStudy (anonymous):

55/63

OpenStudy (anonymous):

\[22/45\div14/25 = 22/45*(25/14) = 55/63\]

OpenStudy (anonymous):

daiel i was lookin for u 2day! geesh

OpenStudy (anonymous):

daniel i was lookin for u 2day! geesh

OpenStudy (anonymous):

thanx tamoy*

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I have to sleep sometime :D

OpenStudy (anonymous):

lol i know rite

OpenStudy (anonymous):

55/63